OpenStudy (anonymous):

Commutative,Associative,Distributive,Symmetric, Transitive.WHAT DO THEY MEAN?i know there to identify number propertys, but what do they mean?

Vocaloid (vocaloid):

quick summary: commutative property: when multiplying or adding numbers, you can change the order without changing the result ex: a + b = b + a, a*b = b*a associative property: when multiplying or adding numbers, you can change the parentheses (order of operations) without changing the result ex. a + (b + c) = (a + b) + c a*(b*c) = (a*b)*c distributive property: uh, this is kinda hard to explain with words, but when you have a quantity multiplied by another quantity involving addition or subtraction, you multiply the outside quantity by each of the terms in the parentheses and simplify ex. a(b+c) = a*b + b*c

Vocaloid (vocaloid):

symmetric property: two quantities on opposite sides of an equals sign can be switched ex. if a = b then b = a transitive property: again, this is kinda hard to explain, maybe an example will make more sense ex. if a = b and b = c, then a = c